# Statistical Inference with Limited Memory: A Survey

### v1

@@ 1 Introduction
Statistical inference is one of the most prominent pillars of modern science and engineering. Generally speaking, a statistical inference problem is the task of reliably estimating some property of interest from random samples, obtained from a partially or completely unknown distribution. Such properties can include for example the mean of the distribution, its higher moments, whether the distribution belongs to a certain class (e.g., is uniform or Gaussian) or is far from that class, the entropy of the distribution, or even the distribution itself. These types of problems arise constantly in many research areas, both theoretical and practical, and across multiple disciplines in engineering, computer science, physics, economics, biology, and many other fields and subfields within.
The extremely broad literature on statistical inference has been concerned almost exclusively with an idealized setting, in which no practical constraints are posed on the inference algorithm or the data collection process. Specifically, two implicit postulates underlying this classical setup are that inference is completely centralized, namely that all the samples are available to a single agent who carries out the inference procedure, and that this agent has essentially unlimited computational power, hence can implement the optimal decision rule. These crucial assumptions, however, often fail to hold in real world settings, where data is either collected in a distributed fashion by multiple remotely-located agents who are subject to stringent communication constraints, or by a single agent who has limited computational capabilities (or both). If the number of collected samples is very small relative to the computational power or the available communication bandwidth, then the computationally-limited / distributed setup can be easily reduced essentially to a computationally unlimited / centralized one with little overhead. This is however not possible when either communication or computation / memory become a bottleneck.
@@
In this survey, we focus on the computational aspect of such problems, and in particular on the limited memory aspect, where an inference algorithm is restricted to store only a small number of bits in memory. The study of learning and estimation under memory constraints can be traced back to the late $1950$s and the problem of the Two Armed Bandit imitated by Robbins [1], and remained an active research area for a few decades. It has then been largely abandoned, but has recently drawn much attention again, with many works addressing different aspects of inference under memory constraints in the last few years. Despite this ongoing effort, there are still gaps in our understanding of the effects that memory limitation has on inference accuracy in various problems, as well as its impact on the required sample size. In this survey, we attempt to provide a snapshot overview of this fast-evolving area, summarizing the main results in the field and the landscape around it. We also delineate a few high level ideas and techniques often used to obtain upper and lower bounds in this type of problems, point out some interesting phenomena and trade-offs, and discuss open problems.
@@ 2.2.1 Simple Hypothesis Testing
which is no other than the error probability of the algorithm. The minimax risk is hence the smallest error probability that can be uniformly guaranteed for all hypotheses in $\mathcal{F}$. A secondary line of work, mostly appearing in the information theory literature, is interested in the decay rate of $\Pr(\pi\neq\hat{\pi})$ as a function of the number of samples $n$, and, particularly, in the optimal error exponent for Neyman–Pearson testing, which is the decay rate of the type II error probability subject to the constraint that the type I error probability is at most some specified constant. For a more symmetric approach, one may consider the Bayesian setting, in which prior probabilities are assigned to both hypotheses and we want to minimize the average error probability, or the minimax setting, in which we want to minimize the maximal error probability. The best achievable error exponent in both cases is the Chernoff information or Chernoff Divergence [9, 10, 11].
@@ 2.2.3 Composite Hypothesis Testing
In this more general problem the family of distributions is a finite disjoint union of sub-families $\mathcal{F}=\cup_{i=1}^{d-1}\mathcal{F}_{i}$, and we want to find out whether the underlying distribution belongs to one of the sub-families. The loss is again the $0/1$ loss thus the expected risk is the minimax error probability. Of a particular interest to us is the distribution property testing variant, on which we elaborate below.
@@ 2.3 Memory Constraints
In the limited-memory framework, an agent observes an i.i.d sequence $X_{1},X_{2},\ldots\in\mathcal{X}$, and at each time point is allowed to update a memory register, or equivalently, a memory state. Throughout the survey, we will use the memory states terminology to refer to a memory limited algorithm, as it is consistent with many of the earlier works and some of the contemporary ones. Specifically, let $\mathsf{FSM}(S,\mathcal{X},\mathcal{Y})$ denote the family of all finite-state machines over the state space $[S]=\{1,2,\ldots,S\}$, with input space $\mathcal{X}$ and output space $\mathcal{Y}$. We will usually write $\mathsf{FSM}(S)$ for short, leaving the input and output spaces implicit. A finite-state machine, or finite-state algorithm, is a pair $(f,g)$ of a state-transition function $f:[S]\times\mathcal{X}\rightarrow[S]$, and a state-decision function $g:[S]\to\mathcal{Y}$. Without loss of generality, we can assume the finite-state machine is initialized to state $1$ at time $n=0$. When fed with the input sequence $X_{1},X_{2},\ldots$, the state of the machine evolves as follows:
@@
where $\mathsf{DFSM}(S)$ is the family of finite-state deterministic estimation algorithms. Let
@@
For any $\delta^{*}\leq\delta$, we can define the $\delta$-sample-memory curve, which is a curve on the $(n,S)$ axis, that represents the minimal number of samples $n$ and the minimal number of states $S$ needed to attain a minimax risk of $\delta$. The curve can be characterized by the function
@@
which gives the minimal number of states $S$ required for the minimax risk to be smaller than $\delta$, or, equivalently, by
@@
which gives the minimal number of samples $n$ required for the minimax risk to be smaller than $\delta$. Note that these two functions are in an inverse relation. The $\delta$-sample-memory curve naturally yields two asymptotic quantities: First, the classical notion of sample complexity is obtained as the minimal number of samples required to attain a risk $\delta$ in the limit of infinite (or large enough) memory:
@@
Note that we define the memory complexity as the number of states the algorithm uses, rather than the number of bits; this is just a matter of convenience, since often writing results using the number of states is more natural and readable, having the same “units” as number of samples. We will nevertheless occasionally use bits for memory size, in which case we will use the term memory complexity in bits.
Characterizing the sample-memory curve in general is a very difficult task, and we are currently unaware of any non-trivial problem for which the $\delta$-sample-memory curve is completely known. Indeed, most of the classical research was focused on only characterizing the two extreme points of $\mathsf{SC}(\delta)$ and $\mathsf{MC}(\delta)$, and only in recent years have other points on the curve been addressed (as we expand upon in later sections). Note that the number of memory bits of an algorithm never exceeds $\mathsf{SC}(\delta)\cdot\mathcal{X}$, as this is the number of bits needed to trivially achieve a risk of $\delta$ by naively saving the first $\mathsf{SC}(\delta)$ samples.
@@ 3 Related Problems
We briefly comment on two models related to inference from i.i.d. samples under memory constraints. The first is the data stream model, which differs from our model in the assumption on the data: in the data stream model no statistical assumptions are made, and the algorithms are designed for the worst case. The second model we consider in this section is the communication-constrained model. In this model, samples are collected by remotely located agents, and the inference bottleneck is imposed by a limited communication budget between the agents. This section is optional and can be skipped.
@@ 3.1 The Data Stream Model
The data stream model characterizes algorithms with small memory size that make passes over the input in order; any item not explicitly stored is inaccessible to the algorithm in the same pass. In many cases the number of passes is limited to one. The figure of merit is taken to be the memory complexity in bits, usually referred to in these works as the space complexity of a problem. The memory complexity in bits of many different streaming algorithms have been extensively studied within theoretical computer science, and some well known examples include estimation of frequency moments of the data stream [13, 14], estimation of Shannon and Rényi entropy of the empirical distribution of the data stream [15, 16, 17], estimation of heavy hitters [18, 19, 20], and estimation of distinct elements [21, 22, 23, 24, 25]. Though some exceptions exist [26, 27], the vast majority of works are concerned with worst case inputs (i.e., adversarially generated data), and thus do not immediately apply in the statistical learning setup, where the data is often assumed to be drawn i.i.d. from some unknown underlying distribution.
@@
In general, the memory consumption of an algorithm designed for a worst case input tends to be significantly larger than the memory needed for the equivalent statistical problem. Consider, for example, the problem of entropy estimation of a stream. To approximate the empirical entropy of a stream of $m$ values up to a multiplicative factor of $(1+\varepsilon)$ using a single pass, one must use at least $\Omega(\varepsilon^{-2}/\log^{2}(\varepsilon))$ memory bits, according to the work of Chakrabarti et al. [17], where $m$ is chosen such that $\varepsilon^{-1}=3\sqrt{m}(\log m+1/2)$. In contrast, when approximating the entropy of a distribution over alphabet $\mathcal{X}=[k]$ up to a multiplicative factor of $(1+\varepsilon)$, only $O(\log(k/\varepsilon))$ memory bits are needed (see [28, 29]). Note that this entails an exponential increase in the number of necessary memory bits w.r.t the dependence on $\varepsilon$.
@@
Another example follows from the problem of approximating the second frequency moment of a sequence, also known as the repeat rate or Gini’s Index of homogeneity. It is well known, as a result of Woodruff [30], that any algorithm that $\varepsilon$ approximates the second moment requires at least $\Omega(1/\varepsilon^{2})$ memory bits, for any $\varepsilon=\Omega(1/\sqrt{m})$, where $m$ is the size of the universe. The statistical variant of the problem is approximating the collision probability of a distribution. Note that this can be done in the following manner: define a collision machine to be a machine that, at each time point, outputs a $1$ if the current sample is the same as the preceding sample and 00 otherwise. This can be implemented using $\log k$ bits, since we are storing the preceding sample to check for collisions, and it results in a Bernoulli process with parameter $p\geq 1/k$. As we show elsewhere in the survey, this can be approximated using $\log(k/\varepsilon)$ bits, resulting in an overall $O(\log(k/\varepsilon))$ memory bits. Thus, again, we have an exponential increase with respect to the loss parameter $\varepsilon$.
@@
There are, however, some interesting rare cases in which the adversarial setting does not cost more memory. A specific example arises from the field of universal prediction with finite-state machines. In a surprising result, Meron and Feder [31] showed that a $K$-state randomized finite-memory predictor of worst case sequences achieves a $\asymp 1/K$ redundancy w.r.t to the least square loss, which is the same redundancy it incurs in the stochastic Bernoulli setting! Another example stems from the problem of $(1+\varepsilon)$ factor distinct elements approximation: It is known, due to the works of Indyk and Woodruff [23] and Kane et al. [25], that the memory complexity in bits of the distinct elements problem, when the input is an $n$-length arbitrary sequence, is $\asymp 1/\varepsilon^{2}+\log n$. Woodruff [32] initiated the study of average-case complexity of the distinct elements problem under certain distributions. Specifically, the case when we observe an $n$-length sequence drawn from a universe of size $m$, where each successive stream item is drawn independently and uniformly at random from an unknown subset of some unknown size $d$. The main contribution of the paper is to show that if $n,d\asymp 1/\varepsilon^{2}$, then the problem requires $\Omega(1/\varepsilon^{2})$ memory bits even in the random data model, thus distinct elements estimation for this family of distributions is as memory-complex as the equivalent adversarial problem.

### v2

@@ 1 Introduction
Statistical inference is one of the most prominent pillars of modern science and engineering. Generally speaking, a statistical inference problem is the task of reliably estimating some property of interest from random samples, obtained from a partially or completely unknown distribution. Such properties can include, for example, the mean of the distribution, its higher moments, whether the distribution belongs to a certain class (e.g., is uniform or Gaussian) or is far from that class, the entropy of the distribution, or even the distribution itself. These types of problems arise constantly in many research areas, both theoretical and practical, and across multiple disciplines in engineering, computer science, physics, economics, biology, and many other fields and subfields within.
The extremely broad literature on statistical inference has been concerned almost exclusively with an idealized setting, in which no practical constraints are posed on the inference algorithm or the data collection process. Specifically, two implicit postulates underlying this classical setup are that inference is completely centralized, namely that all the samples are available to a single agent who carries out the inference procedure, and that this agent has essentially unlimited computational power, hence can implement the optimal decision rule. These crucial assumptions, however, often fail to hold in real world settings, where data is either collected in a distributed fashion by multiple remotely-located agents who are subject to stringent communication constraints, or by a single agent who has limited computational capabilities (or both). [^1] Another common constraint is privacy, and in particular local differential privacy, where the central server is not allowed to learn much about the samples of any individual agent. We refer the interested reader to a survey and a couple of state of the art works on this topic [1, 2, 3]. If the number of collected samples is very small relative to the computational power or the available communication bandwidth, then the computationally-limited / distributed setup can be easily reduced essentially to a computationally unlimited / centralized one with little overhead. However, this is not possible when either communication or computation / memory become a bottleneck.
@@
In this survey, we focus on the computational aspect of such problems, and in particular on the limited memory aspect, where an inference algorithm is restricted to store only a small number of bits in memory. The study of learning and estimation under memory constraints has been initiated in the late 1960s by Cover and Hellman [4, 5] (with a precursor by Robbins [6] on the two armed bandit problem), and remained an active research area for a few decades. It has then been largely abandoned, but has recently drawn much attention again, with many works addressing different aspects of inference under memory constraints in the last few years. Despite this ongoing effort, there are still gaps in our understanding of the effects that memory limitation has on inference accuracy in various problems, as well as its impact on the required sample size. In this survey, we attempt to provide a snapshot overview of this fast-evolving area, summarizing the main results in the field and the landscape around it. We also delineate a few high level ideas and techniques often used to obtain upper and lower bounds in this type of problems, point out some interesting phenomena and trade-offs, and discuss open problems.
Outline. In Section 2 we formally define the general inference setting, present the different types of inference problems we will encounter in the paper, and introduce the setting of inference under finite-state memory constraints. In Section 3, we discuss some assumptions on finite-state algorithms, including time-invariance and randomization. Section 4 is a short digression from the main theme of the paper, where we briefly survey the closely-related problems of inference in the data stream model, and inference under communication constraints. Section 5 introduces the main methodologies used in deriving lower bounds and upper bounds in the memory constrained setting. Sections 6 thru 9 give a thorough overview of the landscape and current status of knowledge in memory constrained hypothesis testing, parameter estimation, distribution testing and distribution property estimation problems. Finally, Section 10 addresses open problems and future research directions.
@@ 2.2.1 Simple Hypothesis Testing
which is precisely the error probability of the algorithm. The minimax risk is hence the smallest error probability that can be uniformly guaranteed for all hypotheses in $\mathcal{F}$. A secondary line of work, mostly appearing in the information theory literature, is interested in the decay rate of $\Pr(\pi\neq\hat{\pi})$ as a function of the number of samples $n$, and, particularly, in the optimal error exponent for Neyman–Pearson testing, which is the decay rate of the type II error probability subject to the constraint that the type I error probability is at most some specified constant. For a more symmetric approach, one may consider the Bayesian setting, in which prior probabilities are assigned to both hypotheses and we want to minimize the average error probability, or the minimax setting, in which we want to minimize the maximal error probability. The best achievable error exponent in both cases is the Chernoff information or Chernoff Divergence [14, 15, 16].
@@ 2.2.3 Composite Hypothesis Testing
In this more general problem the family of distributions is a finite disjoint union of sub-families $\mathcal{F}=\cup_{i=1}^{d-1}\mathcal{F}_{i}$, and we want to find out whether the underlying distribution belongs to one of the sub-families. The loss is again the $0/1$ loss, thus the expected risk is the minimax error probability. Of a particular interest to us is the distribution property testing variant, on which we elaborate below.
@@ 2.3 Memory Constraints
In the limited-memory framework, an agent observes an i.i.d sequence $X_{1},X_{2},\ldots\in\mathcal{X}$, and at each time point is allowed to update a memory register, or, equivalently, a memory state. Throughout the survey, we will use the memory states terminology to refer to a memory limited algorithm, as it is consistent with many of the earlier works and some of the contemporary ones. Specifically, let $\mathsf{FSM}(S,\mathcal{X},\mathcal{Y})$ denote the family of all finite-state machines over the state space $[S]=\{1,2,\ldots,S\}$, with input space $\mathcal{X}$ and output space $\mathcal{Y}$. We will usually write $\mathsf{FSM}(S)$ for short, leaving the input and output spaces implicit. A finite-state machine, or finite-state algorithm, is a pair $(f,g)$ of a state-transition function $f:[S]\times\mathcal{X}\rightarrow[S]$, and a state-decision function $g:[S]\to\mathcal{Y}$. Without loss of generality, we can assume the finite-state machine is initialized to state $1$ at time $n=0$. When fed with the input sequence $X_{1},X_{2},\ldots$, the state of the machine evolves as follows:
@@
where $\mathsf{DFSM}(S)$ is the family of finite-state deterministic estimation algorithms. Let $\delta^{*}\geq 0$ be the smallest obtainable minimax risk, i.e.,
@@
For any $\delta^{*}\leq\delta$, we can define the sample-memory curve at level $\delta$, which is a curve on the $(n,S)$ axis, that represents the minimal number of samples $n$ and the minimal number of states $S$ needed to attain a minimax risk of $\delta$. The curve can be characterized by the function
@@
which gives the minimal number of states $S$ required for the minimax risk to be at most $\delta$, or, equivalently, by
@@
which gives the minimal number of samples $n$ required for the minimax risk to be at most $\delta$. Note that these two functions are in an inverse relation. The sample-memory curve at level $\delta$ naturally yields two asymptotic quantities: First, the classical notion of sample complexity is obtained as the minimal number of samples required to attain a risk $\delta$ in the limit of infinite (or large enough) memory:
@@
Note that we define the memory complexity as the number of states the algorithm uses, rather than the number of bits; this is just a matter of convenience, since often writing results using the number of states is more natural and readable, having the same “units” as number of samples. We will nevertheless occasionally use bits for memory size, in which case we will use the term memory complexity in bits. Note that the memory complexity in bits never exceeds $\mathsf{SC}(\delta)\cdot\log|\mathcal{X}|$, as this is the number of bits needed to trivially achieve a risk of $\delta$ by naively saving the first $\mathsf{SC}(\delta)$ samples.
Characterizing the sample-memory curve in general is a very difficult task, and we are currently unaware of any non-trivial problem for which the sample-memory curve at level $\delta$ is completely known. Indeed, most of the classical research was focused on only characterizing the two extreme points of $\mathsf{SC}(\delta)$ and $\mathsf{MC}(\delta)$, and only in recent years have other points on the curve been addressed (as we expand upon in later sections).
3 Algorithm Types
In this section, we discuss different assumptions that can be made on finite memory algorithms. First, we consider the dependence of the algorithms on time, and argue that it makes sense to restrict the discussion to time-invariant algorithms. We then consider randomized algorithms, i.e. ones that have access to an external source of randomness, and contrast them with deterministic algorithms, where the only randomness comes from the samples.
3.1 Time-Varying vs. Time-Invariant
A finite-memory estimation algorithm $\hat{\pi}$ is said to be time-varying, if either the state-transition function $f$, or the state-decision function $g$ (or both) are allowed to depend the time index $n$. Otherwise, namely if both $f$ and $g$ are fixed for all $n$, the algorithm is called time-invariant. We note that any time-varying algorithm can be trivially reduced to a time-invariant one with an $O(\log{n})$ bits overhead, if the total number of available samples $n$ is finite, by simply keeping a clock. [^2] We do not account here for the complexity of describing the state transition function $f$. In time-varying algorithms, this description may grow as a function of the number of observed samples. Thus, in our framework, a time-varying algorithm implicitly means that we assume an external clock is provided “for free”, i.e., without accounting for the memory required to store it. Since what we care about is the memory budget, it seems less preferable to make this kind of an assumption. Moreover, when the number of samples is unbounded, this assumption can lead to absurd results. For example, one can solve the two-armed Bandit problem with only $2$ states [18], and the binary hypothesis testing problem using just $3$ states and with arbitrarily small error (!) [4, 19]. Finally, as we shall see, some optimal time-invariant algorithms inherently do not keep a clock at all. For all these reasons, in this survey we will restrict our attention to time-invariant algorithms.
3.2 Randomized vs. Deterministic
A randomized algorithm is an algorithm whose state-update and/or state-decision functions depend not only on the samples, but also on some external randomness independent of the samples. A natural question is in what cases randomization even helps? Note that, trivially, there is no loss of generality in assuming that the state-decision function is deterministic; this follows simply since the risk incurred by a random decision rule is some average over risks obtained by deterministic ones, hence one can always pick the risk-minimizing deterministic rule.
@@ 4 Related Problems
We briefly comment on two models related to inference from i.i.d. samples under memory constraints. The first is the data stream model, in which no statistical assumptions are made on the data, and the algorithms are designed for the worst case. The data stream model is hence fundamentally more difficult than the inference setting addressed in this paper. The second setup we consider in this section is the communication-constrained setup. Here, samples are collected by remotely located agents, and the inference bottleneck is imposed by a limited communication budget between the agents. There are several models that capture the limited communication links, some of which are easier then the memory-constrained setup. One such example is the blackboard model, where each agent has access to all the messages sent by previous agents, and each message can take $S$ values. This model is easier than our memory constrained model, in which each agent can choose its message using a finite-state machine whose input is the current observation and the message sent by the previous agent alone. Please note that this section is optional and can be skipped.
@@ 4.1 The Data Stream Model
The data stream model characterizes algorithms with small memory size that make passes over the input in order; any item not explicitly stored is inaccessible to the algorithm in the same pass. In many cases the number of passes is limited to one. The figure of merit is taken to be the memory complexity in bits, usually referred to in these works as the space complexity of a problem. The memory complexity in bits of many different streaming algorithms have been extensively studied within theoretical computer science, and some well known examples include estimation of frequency moments of the data stream [21, 22], estimation of Shannon and Rényi entropy of the empirical distribution of the data stream [23, 24, 25], estimation of heavy hitters [26, 27, 28], and estimation of distinct elements [29, 30, 31, 32, 33]. Though some exceptions exist [34, 35], the vast majority of works are concerned with worst case inputs (i.e., adversarially generated data), and thus lower bounds for the data stream model do not generally apply in the statistical learning setup, where the data is often assumed to be drawn i.i.d. from some unknown underlying distribution.
@@
In general, the data stream model is substantially more difficult than the stochastic setup. To demonstrate this point, we provide a couple of examples for problems where the memory complexity required for evaluating some function of the empirical distribution of a worst case input tends to be significantly larger than the memory complexity for evaluating the same function of the underlying distribution in a stochastic setup. Consider, for example, the problem of entropy estimation of a data stream. To provide an $\varepsilon$ multiplicative approximation of the empirical entropy of a stream of $n$ values in a single pass, one must use at least $\Omega(\varepsilon^{-2}/\log^{2}(\varepsilon))$ memory bits, according to the work of Chakrabarti et al. [25] (the lower bound is proved for the assignment $\varepsilon^{-1}=3\sqrt{n}(\log n+1/2)$). In contrast, when approximating the entropy of a distribution over alphabet $\mathcal{X}=[k]$ with the same multiplicative factor, only $O(\log(k/\varepsilon))$ memory bits are needed (see [36, 37]). This entails an exponential increase in the number of necessary memory bits w.r.t the dependence on $\varepsilon$.
@@
Another example follows from the problem of approximating the second frequency moment of a sequence, also known as the repeat rate or Gini’s Index of homogeneity. Woodruff [38] has shown that any algorithm that $\varepsilon$ approximates the second moment requires at least $\Omega(1/\varepsilon^{2})$ memory bits, for any $\varepsilon=\Omega(1/\sqrt{k})$, where $k$ is the alphabet size. The statistical variant of the problem is approximating the collision probability of a distribution over alphabet $\mathcal{X}=[k]$. A simple algorithm that performs this is the following: define a collision machine to be a machine that, at each time point, outputs $1$ if the current sample is the same as the preceding sample and 00 otherwise. This can be implemented using $\log k$ bits, since we are only storing the preceding sample, and it results in a Bernoulli process with parameter $p\geq 1/k$. As we show elsewhere in the survey, this Bernoulli parameter can be approximated using $O(\log(k/\varepsilon))$ bits, resulting in an overall $O(\log(k/\varepsilon))$ memory bits. Thus, again, we have an exponential increase with respect to the loss parameter $\varepsilon$.
@@
There are, however, some interesting rare cases in which the adversarial setting does not cost more memory than the stochastic one. A specific example arises from the field of universal prediction with finite-state machines. In a surprising result, Meron and Feder [39] showed that a $K$-state randomized finite-memory predictor for worst case sequences achieves a $\asymp 1/K$ redundancy w.r.t to the least square loss, which is the same redundancy it incurs in the stochastic Bernoulli setting. Another example stems from the problem of distinct elements approximation: It is known, due to the works of Indyk and Woodruff [31] and Kane et al. [33], that the memory complexity in bits of the distinct elements problem is $\asymp 1/\varepsilon^{2}+\log n$, when the input length is $n$ and $\varepsilon$ is the approximation parameter. Woodruff [40] initiated the study of average-case complexity of the distinct elements problem under certain distributions and, in particular, considered a stochastic setup in which we observe an $n$-length random sequence drawn from an alphabet size $k$, where each element is drawn uniformly and independently from an unknown subset of some unknown size $d$. The main contribution of the paper is to show that if $n,d\asymp 1/\varepsilon^{2}$, then the problem requires $\Omega(1/\varepsilon^{2})$ memory bits. Thus, even in the random data model, distinct elements estimation can be as memory-complex as its data stream variant.